OpenStudy (anonymous):

You are fencing a rectangular puppy kennel with 25 ft of fence. The side of the kennel against your house does not need a fence. This side is 9 ft long. Find the dimensions of the kennel.

OpenStudy (anonymous):

Perimeter = 2l + 2w 25ft + 9ft = 2(9) + 2w 34ft = 18 + 2w 16ft = 2w w = 8 ft answer : length = 9 ft width = 8ft